OpenStudy (anonymous):

can someone help me they put trig in my algebra 2 class i really need help

jimthompson5910 (jim_thompson5910):

what is your question?

OpenStudy (anonymous):

actually i have a bunch of questions lolo i dont know what any of this is

jimthompson5910 (jim_thompson5910):

ok i'll help you out with a few, but I won't do all of them

OpenStudy (anonymous):

in triangle ABC, Angle C is a right angle. Find the remaining sides and angles. Round your answer to the nearest 10th a = 3 c =19

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

let's find b first a^2 + b^2 = c^2 ... start with the pythagorean theorem 3^2 + b^2 = 19^2 9 + b^2 = 361 keep going to solve for b

OpenStudy (anonymous):

umm im not sure what to do next

jimthompson5910 (jim_thompson5910):

b^2 = 361 - 9 b^2 = 352 b = ???

OpenStudy (anonymous):

subract 19^2 from 3^2?

jimthompson5910 (jim_thompson5910):

you mean 3^2 from 19^2

OpenStudy (anonymous):

o yes i got 352

jimthompson5910 (jim_thompson5910):

then take the square root of that to get b

OpenStudy (anonymous):

idk how to do that

jimthompson5910 (jim_thompson5910):

do you have a calc?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{352} = 18.76166304 ?????\]

jimthompson5910 (jim_thompson5910):

round that to the nearest tenth

OpenStudy (anonymous):

im not sure what nearest 10th is i never learned that

jimthompson5910 (jim_thompson5910):

you never learned rounding?

OpenStudy (anonymous):

no sir

jimthompson5910 (jim_thompson5910):

that's quite odd, but ok 18.76166304 the tenth place is the 7 the digit after that is a 6 (greater than 5), so we drop everything after the 7 and bump the 7 up to 8 so after rounding 18.76166304 to the nearest tenth, we get 18.8

jimthompson5910 (jim_thompson5910):